OpenStudy (anonymous):

What is the area of a sector with radius 6" and measure of arc equal to 120°?

hero (hero):

\[\frac{120}{360} = \frac{x}{ \pi (6)^2}\]

hero (hero):

Solve for x

hero (hero):

In general use this formula: \[\frac{\theta}{360} = \frac{x}{2 \pi r}\] where x = area of sector

OpenStudy (anonymous):

Area would be \[(\pi*r^{2}) * 120/360\] = \[(\pi*6^2)/3 \approx 37.7 units\]